OpenStudy (anonymous):

hey.Help me here please. sin (x-pi/4)-1=0

OpenStudy (ikram002p):

\(\large \sin (x-\frac{\pi}{4})-1=0\) u wanna find x right ?

OpenStudy (ikram002p):

let \(x-\frac{\pi}{4}=y\) \(\large \sin (y)-1=0\) \(\large \sin (y)=1\)

OpenStudy (ikram002p):

now wich \(y\) value gives u \(\large \sin (y)=1\) \(\large y=\frac{n \pi}{2}\) right ? cuz \(\large \sin (\frac {\pi}{2})=1\)

OpenStudy (ikram002p):

got it untill nw ?

OpenStudy (anonymous):

thanks

OpenStudy (ikram002p):

ur wlc , but still u need to find x :)

OpenStudy (anonymous):

ok

OpenStudy (ikram002p):

:) ok

OpenStudy (anonymous):

\[y=\frac{ \pi }{ 2 }+2\pi k\] \[k=0,\pm1,\pm2,......\]